For the first order gas phase reaction of an organic compound at 130 0C at                     Ea is 108.4 kJmol-1and rate constant is 9.12x10-4 s-1. Calculate the standard entropy ΔS# of this reaction using eyring equation.


1
Expert's answer
2021-07-08T04:54:44-0400

K = KbT/h e∆G/RT


K = KbT/h e∆S/R e-Ea/RT


9.12x10-4 = 8.617×10−5 x 403/6.63 × 10-34 x e∆S/8.314 x e-108.4/8.314 x403

e∆S/8.314 = 1.75 x10-35

∆S/8.314 = -80.00

∆S= -665.15J/k